OpenStudy (anonymous):

Find the polynomial function with roots 1, 7, and -3 of multiplicity 2.

OpenStudy (campbell_st):

isn't it simply y = a(x - 1)(x + 7)(x + 3)^2

OpenStudy (anonymous):

I don't know. This is my first time ever having to do math like this.

OpenStudy (campbell_st):

ok... so roots are where the curve cuts the x axis if a root is x = 3 then in a factored polynomial it is written ( x -3) this may help to explain the opposite signs multiplicity of 2 means the root is repeated twice (x + 3)(x+3) = (x + 3)^2

OpenStudy (anonymous):

Okay I understand.